# FinVet: A Collaborative Framework of RAG and External Fact-Checking Agents for Financial Misinformation Detection

### v1

@@ Abstract
Financial markets face growing threats from misinformation that can trigger billions in losses in minutes. Most existing approaches lack transparency in their decision-making and provide limited attribution to credible sources. We introduce FinVet, a novel multi-agent framework that integrates two Retrieval-Augmented Generation (RAG) pipelines with external fact-checking through a confidence-weighted voting mechanism. FinVet employs adaptive three-tier processing that dynamically adjusts verification strategies based on retrieval confidence, from direct metadata extraction to hybrid reasoning to full model-based analysis. Unlike existing methods, FinVet provides evidence-backed verdicts, source attribution, confidence scores, and explicit uncertainty flags when evidence is insufficient. Experimental evaluation on the FinFact dataset shows that FinVet achieves an F1 score of 0.85, which is a 10.4% improvement over the best individual pipeline (fact-check pipeline) and 37% improvement over standalone RAG approaches.
@@ I Introduction
Financial markets are increasingly susceptible to misinformation, where a single false claim can trigger billions of dollars in losses within minutes. For instance, a 2013 tweet from a hacked Associated Press account falsely reporting a White House explosion “wiped out \$130 billion from the stock market” [1]. More recently, in May 2023, a viral fake image of a Pentagon explosion caused the Dow Jones Industrial Average to drop 80 basis points (bps) and the S&P 500 to fall by 26 bps [2]. These incidents underscore the urgent need for a robust and transparent framework capable of detecting and mitigating financial misinformation. Financial misinformation is defined as false or misleading information related to economic or financial matters that can influence investor decisions, market behavior, or regulatory outcomes [3]. It can also include fraudulent schemes aimed at defrauding individuals or institutions.
The advent of generative AI has triggered a wave of AI-driven tools, escalating misinformation, disrupting financial markets, and fueling a surge in fraud. Malicious actors now exploit generative AI to create deepfake videos, synthetic voices, and fraudulent documents—enabling large-scale deception that challenges financial institutions’ detection capabilities. Social media platforms, central to the rapid dissemination of misinformation, exacerbate this growing threat. According to Deloitte [4], AI-enabled fraud losses in the United States are projected to increase from \$12.3 billion in 2023 to \$40 billion by 2027.
Identifying and mitigating financial misinformation is crucial for financial institutions, investors, and regulators. Left unchecked, misinformation can manipulate stock prices, mislead investors, and destabilize markets. To this end, this paper introduces FinVet, a novel multi-agent framework designed to address the aforementioned challenges through the integration of two Retrieval-Augmented Generation (RAG) [5] pipelines and a Fact-Check Pipeline. The RAG components utilize an external knowledge source, such as a vector store, which can be tailored to any domain-specific dataset. The Fact-Check Pipeline combines an external fact-checking source with a fallback mechanism leveraging a Large Language Model (LLM) when direct evidence is unavailable. The outputs of these independent pipelines are integrated via a confidence-based voting mechanism to deliver a final verdict. This integrated approach provides a robust and adaptable misinformation detection system, offering improved accuracy and explainability. In addition to flagging misinformation, FinVet presents evidence-backed justifications, clearly cites source origins, and assigns a confidence score quantifying the system’s certainty in its verdict.
@@ II-A Deep Learning-Based Techniques
Dmonte et al. [6] instruction-tune LLaMA-3.1-8B using claim–justification pairs and also explore a prompting-based few-shot approach. Similarly, Purbey et al. [7] use a sequential fine-tuning strategy—first training LLMs for classification and then for joint explanation generation—to optimize end-to-end performance on the financial misinformation detection task.
Zhang and Liu [8] focus on identifying financial news in the Chinese market with a deep learning approach. They propose a dual sub-network model, employing BERT-Chinese-wwm embeddings [9][10] to generate rich word representations, which feed into three parallel Convolutional Neural Networks (CNNs) with varying kernel sizes to extract multi-granular semantic features from article content.
Kamal, Mohankumar, and Singh[11] tackle financial misinformation with their Fin-MisID model. They leverage RoBERTa [12] for contextual representations, feeding them into multi-channel networks of parallel Convolutional Neural Networks (CNNs) with varying kernel sizes, Bidirectional Gated Recurrent Units (BiGRU) for sequential patterns, and attention mechanisms to emphasize critical tokens.
Nasir, Khan, and Varlamis[13] propose a hybrid deep learning architecture combining CNN and Long Short-Term Memory networks (LSTMs) for fake news detection. CNNs are used to extract local textual patterns and n-gram features, which are then passed to LSTMs to capture long-range dependencies in the sequence. The model is trained end-to-end, enabling joint learning of spatial and temporal features. This approach leverages the complementary strengths of CNNs and RNNs for improved representation of news content.
X. Zhang, Q. Du, and Z. Zhang[14] design a theory-driven machine learning system for financial disinformation detection, grounded in Truth-Default Theory (TDT). Their framework operationalizes five key deception cues—communication context and motive, sender demeanor, third-party information, coherence, and correspondence—into quantifiable features. These features are extracted from financial news articles, propagation patterns across platforms, author metadata, and firm-level financial indicators. The system integrates these heterogeneous signals into a supervised learning pipeline, offering interpretable and domain-sensitive disinformation detection.
Zhi et al.[15] propose a multi-fact CNN-LSTM model that integrates multiple dimensions of information for financial fake news detection. The model combines news content, market data, user comments, and source credibility into a unified framework. Character-level CNNs are employed to extract granular textual features, which are then encoded through LSTM layers to capture temporal dependencies. An attention mechanism models the interaction between news content and user comments, and all components are fused through weighted aggregation for the final prediction.
@@ II-B LLM-based Approaches
Large Language Models (LLMs) have recently been explored for misinformation detection. Leite, Razuvayevskaya, Bontcheva, and Scarton [16] analyze whether LLMs can generate weak labels based on 18 credibility signals, which are then aggregated using weak supervision to predict content veracity. Their approach demonstrates that zero-shot LLM-based credibility labeling, combined with signal aggregation, can outperform traditional classifiers without requiring ground-truth labels.
Liu et al.[17] developed an instruction-tuning dataset, the Financial Misinformation Detection Instruction Dataset (FMDID), by combining two existing datasets: FinFact [3] and FinGuard [18]. Using their developed dataset, the authors fine-tuned LLaMa2-chat-7b [19] and Llama-3.1-8B-Instruct  [^1] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct. They compared their model’s performance against baseline models, including Pre-trained Language Models (PLMs) like BERT and RoBERTa, as well as LLMs such as LLaMA2, LLaMA3.1, and ChatGPT [20].
Wan et al. [21] propose DELL, a multi-stage misinformation detection framework that strategically integrates LLMs to improve veracity assessment. DELL begins by generating synthetic user-news interaction networks through various LLM-generated news reactions. It uses six explainable proxy tasks, such as sentiment, stance, and framing to enrich contextual understanding with LLM-generated explanations. Finally, it employs an LLM-based expert ensemble, where task-specific predictions and confidence scores are selectively aggregated to produce calibrated, rationale-backed misinformation judgments.
@@ IV Methodology
This research proposes FinVet shown in Figure 1 and elaborated in Algorithm 1. FinVet is a novel framework that employs a multi-pipeline architecture for financial claim verification. The framework consists of four primary components: (A) Data Processing and Vector Storage, (B) Claim Verification Pipelines, (C) Results Normalization, and (D) Verdict Integration and Reporting. These components are described in the following sections.
@@ IV-B1 RAG Verification Pipelines
The RAG Verification Pipelines implement a retrieval-augmented generation approach that grounds verification in contextual knowledge. Each pipeline retrieves relevant financial information from the vector store and uses this context in its verification process. FinVet employs two distinct RAG pipelines that are implemented using different underlying langugage models. The framework is inherently extensible, allowing for the integration of additional RAG pipelines to further enhance the verification process. This modular design enables the incorporation of specialized verification pathways tailored to particular financial domains or claim types, such as market predictions, earnings reports, or macroeconomic indicators. Moreover, while the current RAG implementations use a shared vector store, the architecture supports the integration of multiple domain-specific knowledge bases that can be used for different financial sectors, regulatory frameworks, or historical time periods, enabling the system to triangulate evidence using multiple sources.
@@
  • Case 2: For moderately relevant evidence (0.4 <=<= Similarity Score << 0.6), a hybrid approach is used by combining the retrieved context with model reasoning, providing both the retrieved information and claim to the foundation model with explicit instructions to evaluate the claim using the provided context. The evidence is generated by the foundation model and the source attribution maintains the original retrieved sources when the model validates their relevance; if the model determines the retrieved information is insufficient, it indicates ”Parametric Knowledge” as the source. Confidence is calculated as the average of the retrieval similarity and the model’s self-reported confidence, balancing evidence relevance with model certainty;
  • Case 3: When no strongly relevant evidence is available (similarity << 0.4), the system falls back to a role-based prompting that instructs the foundation model to analyze the claim from multiple expert perspectives (Financial Analyst, Political Misinformation Specialist, Government Policy Analyst, and Investigative Journalist). In this tier, evidence is entirely generated by the foundation model’s reasoning process and explicitly identified as model-generated content. Source attribution is consistently marked as ”Parametric Knowledge” to maintain transparent distinction between retrieved and generated content. Confidence is based entirely on calibrated assessment of the foundation model.
@@ IV-B2 Fact-Check Pipeline
The Fact-Check Pipeline provides a complementary verification approach that augments the context-based RAG methodology. This pipeline operates through two distinct verification stages:
  • External Verification Component: The system queries an external fact-checking source to identify previously analyzed claims. When matches are found, the system extracts evidence from the fact-check response, prioritizing textual ratings and publisher conclusions. Source attribution is preserved directly via the URL to the original fact-checked source. For cases where a match is available in the external source, the system assigns the maximum confidence (1.0) to the verdict.
  • LLM Analyzer: When external verification yields no results, the system uses role-based prompting to an LLM. Evidence is generated through the structured analysis of the claim by the model, and sources are explicitly labeled as ”Parametric Knowledge”. Confidence scores are extracted from the model’s self-reported confidence rating.
@@ IV-C Results Normalization
At this stage, FinVet has the verification results of a claim using three independent pipelines: two RAG pipelines and an external Fact-Check Pipeline. This component transforms these diverse outputs into a consistent format. It normalizes labels to lower case (”true”, ”false” or ”nei”) and standardizes confidence score formatting which is crucial for the next and final component.
@@ IV-D Verdict Integration and Reporting
    If all pipelines return zero confidence scores, the system defaults to a ”NEI” verdict, explicitly acknowledging insufficient evidence rather than making an unwarranted classification.
@@ V-A1 Vector Embedding and Storage
FinVet was implemented using the text embedding model all-MiniLM-L6-v2 [22][23]. This sentence transformer model generates compact 384-dimensional embeddings that effectively capture semantic relationships between the concepts of a claim. The Facebook AI Similarity Search (FAISS) library with InVerted File with Flat storage (IVFFlat) architecture [24] was used to index and retrieve the embeddings, enabling fast similarity search in large volumes of financial statements.
@@ V-A2 RAG Pipeline Implementation
  • RAG Pipeline-1 was implemented using the LLaMA-3.3-70B-Instruct model [^2] https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct [25], accessed through the Hugging Face API. This transformer-based model was selected for its strong general knowledge and reasoning capabilities.
  • RAG Pipeline-2 was implemented using the Mixtral-8x7B-Instruct-v0.1 model [^3] https://huggingface.co/mistralai/Mixtral-8x7B-Instruct-v0.1 [26], accessed through the Hugging Face API. This model was selected for its complementary reasoning patterns through its Mixture-Of-Experts (MOE) architecture.
@@ V-A3 Fact-Check Pipeline
  • The external verification component was interfaced with the Google Fact Check Tools API [27].
@@ V-B Dataset and Evaluation Methodology
FinVet was evaluated using the FinFact dataset [3], a comprehensive collection of financial claims with expert-verified labels and supporting evidence. The dataset was partitioned into 85% training and 15% testing datasets. The training dataset was used to populate the vector store, while the remaining test dataset was used for performance evaluation purposes. The approaches are evaluated using standard metrics: accuracy (Acc.), precision (Prec.), recall (Rec.), and F1-score (F1).

### v2

@@ Abstract
Financial markets face growing threats from misinformation that can trigger billions of dollars in losses within minutes. Most existing approaches lack transparency in their decision-making and provide limited attribution to credible sources. We introduce FinVet, a novel multi-agent framework that integrates two Retrieval-Augmented Generation (RAG) pipelines with external fact-checking through a confidence-weighted voting mechanism. FinVet employs adaptive three-tier processing that dynamically adjusts verification strategies based on retrieval confidence, from direct metadata extraction to hybrid reasoning to full model-based analysis. Unlike existing methods, FinVet provides evidence-backed verdicts, source attribution, confidence scores, and explicit uncertainty flags when evidence is insufficient. Experimental evaluation on the FinFact dataset shows that FinVet achieves an F1 score of 0.85, which is a 10.4% improvement over the best individual pipeline (Fact-Check pipeline) and 37% improvement over standalone RAG approaches.
@@ I Introduction
Financial markets are increasingly susceptible to misinformation, in which a single false claim can trigger billions of dollars in losses within minutes. For instance, a 2013 tweet from a hacked Associated Press account falsely reporting a White House explosion “wiped out \$130 billion from the stock market" [Qian2018]. More recently, in May 2023, a viral fake image of a Pentagon explosion caused the Dow Jones Industrial Average to drop 80 basis points (bps) and the S&P 500 to fall by 26 bps [Parsan2024]. These incidents underscore the urgent need for a robust and transparent framework capable of detecting and mitigating financial misinformation. Financial misinformation is defined as false or misleading information related to economic or financial matters that can influence investor decisions, market behavior, or regulatory outcomes [Rangapur2024]. It can also include fraudulent schemes aimed at defrauding individuals or institutions.
The advent of generative AI has triggered a wave of AI-driven tools, escalating misinformation, disrupting financial markets, and fueling a surge in fraud. Malicious actors now exploit generative AI to create deepfake videos, synthetic voices, and fraudulent documents—enabling large-scale deception that challenges financial institutions’ detection capabilities. Social media platforms, central to the rapid dissemination of misinformation, exacerbate this growing threat. According to Deloitte [Deloitte2023], AI-enabled fraud losses in the United States are projected to increase from \$12.3 billion in 2023 to \$40 billion by 2027.
Identifying and mitigating financial misinformation is crucial for financial institutions, investors, and regulators. Left unchecked, misinformation can manipulate stock prices, mislead investors, and destabilize markets. To this end, this paper introduces FinVet, a novel multi-agent framework designed to address the aforementioned challenges through the integration of two Retrieval-Augmented Generation (RAG) [lewis2020] pipelines and a Fact-Check Pipeline. The RAG components utilize an external knowledge source, such as a vector store, which can be tailored to any domain-specific dataset. The Fact-Check Pipeline combines an external fact-checking source with a fallback mechanism leveraging a Large Language Model (LLM) when direct evidence is unavailable. The outputs of these independent pipelines are integrated via a confidence-based voting mechanism to deliver a final verdict. This integrated approach provides a robust and adaptable misinformation detection system, offering improved accuracy and explainability. In addition to flagging misinformation, FinVet presents evidence-backed justifications, clearly cites source origins, and assigns a confidence score quantifying the system’s certainty in its verdict.
@@ II-A Deep Learning-Based Techniques
Dmonte et al. [dmonte2025gmu] instruction-tune LLaMA-3.1-8B using claim–justification pairs and also explore a prompting-based few-shot approach. Similarly, Purbey et al. [purbey2025sharedtasks] use a sequential fine-tuning strategy: first, training LLMs for classification, then for joint explanation generation to optimize end-to-end performance on the financial misinformation detection task.
Zhang and Liu [Zhang2023] focus on identifying financial news in the Chinese market with a deep learning approach. They propose a dual sub-network model, employing BERT-Chinese-wwm embeddings [Devlin2018],[cui2021] to generate rich word representations, which feed into three parallel Convolutional Neural Networks (CNNs) with varying kernel sizes to extract multi-granular semantic features from article content.
Kamal, Mohankumar, and Singh[Kamal2023] tackle financial misinformation with their Fin-MisID model. They leverage RoBERTa [Liu2019] for contextual representations, feeding them into multi-channel networks of parallel Convolutional Neural Networks (CNNs) with varying kernel sizes, Bidirectional Gated Recurrent Units (BiGRU) for sequential patterns, and attention mechanisms to emphasize critical tokens.
Nasir, Khan, and Varlamis[Nasir2021] propose a hybrid deep learning architecture combining CNN and Long Short-Term Memory networks (LSTMs) for fake news detection. CNNs are used to extract local textual patterns and n-gram features, which are then passed to LSTMs to capture long-range dependencies in the sequence. The model is trained end-to-end, enabling joint learning of spatial and temporal features. This approach leverages the complementary strengths of CNNs and RNNs for improved representation of news content.
X. Zhang, Q. Du, and Z. Zhang[zhang2022] design a theory-driven machine learning system for financial disinformation detection, grounded in Truth-Default Theory (TDT). Their framework operationalizes five key deception cues—communication context and motive, sender demeanor, third-party information, coherence, and correspondence—into quantifiable features. These features are extracted from financial news articles, propagation patterns across platforms, author metadata, and firm-level financial indicators. The system integrates these heterogeneous signals into a supervised learning pipeline, offering interpretable and domain-sensitive disinformation detection.
Zhi et al.[zhi2021financial] propose a multi-fact CNN-LSTM model that integrates multiple dimensions of information for financial fake news detection. The model combines news content, market data, user comments, and source credibility into a unified framework. Character-level CNNs are employed to extract granular textual features, which are then encoded through LSTM layers to capture temporal dependencies. An attention mechanism models the interaction between news content and user comments, and all components are fused through weighted aggregation for the final prediction.
@@ II-B LLM-based Approaches
Large Language Models (LLMs) have recently been explored for misinformation detection. Leite, Razuvayevskaya, Bontcheva, and Scarton [leite2024] analyze whether LLMs can generate weak labels based on 18 credibility signals, which are then aggregated using weak supervision to predict content veracity. Their approach demonstrates that zero-shot LLM-based credibility labeling, combined with signal aggregation, can outperform traditional classifiers without requiring ground-truth labels.
Liu et al.[Liu2024] developed an instruction-tuning dataset, the Financial Misinformation Detection Instruction Dataset (FMDID), by combining two existing datasets: FinFact [Rangapur2024] and FinGuard [Martin2021]. Using their developed dataset, the authors fine-tuned LLaMa2-chat-7b [Touvron2023] and Llama-3.1-8B-Instruct  [^1] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct. They compared their model’s performance against baseline models, including Pre-trained Language Models (PLMs) like BERT and RoBERTa, as well as LLMs such as LLaMA2, LLaMA3.1, and ChatGPT [OpenAI2024].
Wan et al. [wan2024] propose DELL, a multi-stage misinformation detection framework that strategically integrates LLMs to improve veracity assessment. DELL begins by generating synthetic user-news interaction networks through various LLM-generated news reactions. It uses six explainable proxy tasks, such as sentiment, stance, and framing to enrich contextual understanding with LLM-generated explanations. Finally, it employs an LLM-based expert ensemble, where task-specific predictions and confidence scores are selectively aggregated to produce calibrated, rationale-backed misinformation judgments.
@@ IV Methodology
This research proposes FinVet shown in Figure 1, and elaborated in Algorithm 1. FinVet is a novel framework that employs a multi-pipeline architecture for financial claim verification. The framework consists of four primary components: (A) Data Processing and Vector Storage, (B) Claim Verification Pipelines, (C) Results Normalization, and (D) Verdict Integration and Reporting. These components are described in the following sections.
@@ IV-B1 RAG Verification Pipelines
The RAG Verification Pipelines implement a retrieval-augmented generation approach that grounds verification in contextual knowledge. Each pipeline retrieves relevant financial information from the vector store and uses this context in its verification process. FinVet employs two distinct RAG pipelines that are implemented using different underlying language models. The framework is inherently extensible, allowing for the integration of additional RAG pipelines to further enhance the verification process. This modular design enables the incorporation of specialized verification pathways tailored to particular financial domains or claim types, such as market predictions, earnings reports, or macroeconomic indicators. Moreover, while the current RAG implementations use a shared vector store, the architecture supports the integration of multiple domain-specific knowledge bases that can be used for different financial sectors, regulatory frameworks, or historical time periods, enabling the system to triangulate evidence using multiple sources.
@@
  • Case 2: For moderately relevant evidence (0.4 <=<= Similarity Score << 0.6), a hybrid approach is used by combining the retrieved context with model reasoning, providing both the retrieved information and claim to the foundation model with explicit instructions to evaluate the claim using the provided context. The evidence is generated by the foundation model and the source attribution maintains the original retrieved sources when the model validates their relevance; if the model determines the retrieved information is insufficient, it indicates "Parametric Knowledge" as the source. Confidence is calculated as the average of the retrieval similarity and the model’s self-reported confidence, balancing evidence relevance with model certainty;
  • Case 3: When no strongly relevant evidence is available (similarity << 0.4), the system falls back to a role-based prompting that instructs the foundation model to analyze the claim from multiple expert perspectives (Financial Analyst, Political Misinformation Specialist, Government Policy Analyst, and Investigative Journalist). In this tier, evidence is entirely generated by the foundation model’s reasoning process and explicitly identified as model-generated content. Source attribution is consistently marked as "Parametric Knowledge" to maintain transparent distinction between retrieved and generated content. Confidence is based entirely on calibrated assessment of the foundation model.
@@ IV-B2 Fact-Check Pipeline
The Fact-Check Pipeline provides a modular verification layer that complements the RAG-based methodology by linking FinVet to externally verified, authoritative information sources. It can interface with any structured fact-checking repository, whether academic, governmental, or institutional. The pipeline operates through two distinct stages:
  • External Verification Component: This component queries externally maintained fact-checking repositories to determine whether the input claim has already been verified. If a matching record exists, the system retrieves the publisher’s verdict, explanatory rationale, and source reference, then maps these elements into FinVet’s standardized output format comprising verdict, evidence, and source. Verified claims obtained from external repositories are treated as authoritative and assigned a fixed confidence value of 1.0. In practice, this component can connect to publicly available repositories such as Google Fact Check Tools, PolitiFact, Snopes, or Reuters Fact Check, while the architecture also supports integration with other domain-specific fact-checking APIs.
  • LLM Analyzer: When no verified record is found externally, the system employs role-based prompting to an LLM to analyze the claim and generate verdict and supporting evidence. The source of the verdict from this stage is explicitly labeled ’Parametric Knowledge’, distinguishing model-generated reasoning from externally verified information. The confidence value is also derived from the LLM’s internal scoring or self-assessed certainty estimate.
To limit potential error propagation, the prompting template constrains the model to cite only information derived from the input claim, to flag ambiguity or insufficient context with a “Not Enough Information (NEI)” label instead of producing unsupported conclusions, and to perform basic internal consistency checks such as temporal and numerical validation. These measures ensure that the reasoning process remains bounded, transparent, and traceable within FinVet’s verification workflow.
@@ IV-C Results Normalization
At this stage, FinVet has the verification results of a claim using three independent pipelines: two RAG pipelines and an external Fact-Check Pipeline. This component transforms these diverse outputs into a consistent format. It normalizes labels to lower case ("true", "false" or "nei") and standardizes confidence score formatting which is crucial for the next and final component.
@@ IV-D Verdict Integration and Reporting
    If all pipelines return zero confidence scores, the system defaults to a "NEI" verdict, explicitly acknowledging insufficient evidence rather than making an unwarranted classification.
@@ V-A1 Vector Embedding and Storage
FinVet was implemented using the text embedding model all-MiniLM-L6-v2 [wang2020][reimers2019]. This sentence transformer model generates compact 384-dimensional embeddings that effectively capture semantic relationships between the concepts of a claim. The Facebook AI Similarity Search (FAISS) library with InVerted File with Flat storage (IVFFlat) architecture [jegou2022] was used to index and retrieve the embeddings, enabling fast similarity search in large volumes of financial statements.
@@ V-A2 RAG Pipeline Implementation
  • RAG Pipeline-1 was implemented using the LLaMA-3.3-70B-Instruct model [^2] https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct [grattafiori2024], accessed through the Hugging Face API. This transformer-based model was selected for its strong general knowledge and reasoning capabilities.
  • RAG Pipeline-2 was implemented using the Mixtral-8x7B-Instruct-v0.1 model [^3] https://huggingface.co/mistralai/Mixtral-8x7B-Instruct-v0.1 [jiang2024], accessed through the Hugging Face API. This model was selected for its complementary reasoning patterns through its Mixture-Of-Experts (MOE) architecture.
@@ V-A3 Fact-Check Pipeline
  • The external verification component was interfaced with the Google Fact Check Tools API [googlefactcheck2025].
@@
The LLMs in this study were configured with a temperature of 0.10 to ensure a more deterministic output and max_new_tokens was set to 300 to balance response completeness with computational efficiency.
@@ V-B Dataset and Evaluation Methodology
FinVet was evaluated using the FinFact dataset [Rangapur2024], a comprehensive collection of financial claims with expert-verified labels and supporting evidence.
TABLE I: Label Distribution in the FinFact Dataset
Label Samples Percentage
False 1,492 44.29%
True 1,275 37.85%
Not Enough Information (NEI) 602 17.87%
Total 3,369 100.00%
The dataset exhibits a moderate class imbalance, as summarized in Table I. To preserve representative class proportions across training and testing phases, stratified sampling was applied, allocating 85% of the data for training and 15% for testing. The training subset was utilized to populate the vector store, whereas the test subset was reserved exclusively for performance evaluation. Model performance was assessed using standard metrics, including accuracy (Acc.), precision (Prec.), recall (Rec.), and F1-score (F1). To address the effect of class imbalance, the F1-score was adopted as the primary evaluation metric in all comparative analyses.